OpenStudy (anonymous):

What is the slope and y-intercept of the graph of 3x + 2y = 12?

OpenStudy (anonymous):

First you have to isolate y. Put it in y=mx+b form

OpenStudy (anonymous):

how do u do that

OpenStudy (anonymous):

3x+2y=12 2y=12-3x y=12/2-3x/2 y=-3/2x+6

OpenStudy (anonymous):

So now that that is done just pull the information you need. Slope is -3/2 and y intercept is 6

OpenStudy (anonymous):

Your welcome

OpenStudy (anonymous):

thanks